OpenStudy (xximhisgirlxx):

Order from least to greatest 1/2 3/2 2/5 1/3 I'm not good at fractions :/ @jim_thompson5910

OpenStudy (xximhisgirlxx):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

convert each fraction to a decimal and sort that way

jimthompson5910 (jim_thompson5910):

so 1/2 = 0.5 3/2 = 1.5 etc etc

OpenStudy (xximhisgirlxx):

how do you convert?

jimthompson5910 (jim_thompson5910):

using a calculator

jimthompson5910 (jim_thompson5910):

divide 1/2 to get 0.5 using a calculator etc etc

OpenStudy (xximhisgirlxx):

so it's 0.5, 1.5, 0.4, 0.3 ? Now from least to greatest ; 0.3, 0.4, 0.5, 1.5 ?

jimthompson5910 (jim_thompson5910):

1/3 = 0.3333... the 3's repeat forever, but you get the idea

jimthompson5910 (jim_thompson5910):

because the order is 0.33333, 0.4, 0.5, 1.5 this means the order is 1/3, 2/5, 1/2, 3/2

jimthompson5910 (jim_thompson5910):

you replace each decimal number with its fraction counterpart (eg: replace 0.4 with 2/5)